OpenStudy (anonymous):

f(x)=1/x-2+1. Find f(3). Options: 6/5 5/2 2 -3

OpenStudy (saifoo.khan):

\[\frac{1}{x} -2+1 ???\]

OpenStudy (saifoo.khan):

like this?

OpenStudy (anonymous):

1 --- + 1 x-2

OpenStudy (saifoo.khan):

Alright.

OpenStudy (saifoo.khan):

write 3 instead of x.

OpenStudy (anonymous):

Yes but what would the answer be out of the options given? :/

OpenStudy (saifoo.khan):

\[\frac{1}{3-2} +1\] \[\frac{1}{1}+1\] 1+1 2